The value of your stock investment decreased by 32% after a stock market crash. What percentage increase in value would the stocks have to rise in order to return to the value they were before the stock market crash? Round your answer to the nearest tenth of a percent.
Answer #1

Let the initial value of investment be 'K'

Value of investment after crash = K x (1 - 0.32) = 0.68K

To restore the value to K, the rise required is K/0.68K = 1/0.68 = 1.47

The current value have to increase by a factor of 1.47, that is by (1.471 - 1) x 100 = 47.1%